OpenStudy (anonymous):

hey there i'm kinda new here but here it goes how can i solve systems of equations

OpenStudy (anonymous):

there r many methods i personally like to use matrices and gaussian elimination

OpenStudy (anonymous):

can yo do it without a grap

OpenStudy (anonymous):

*graph*

OpenStudy (amistre64):

GRAPHS ARE RATHER IMPRECISE ... im gonna break that pinky

OpenStudy (amistre64):

many systems are not graphable to begin with

OpenStudy (anonymous):

sure for example suppose u have the system\[3x+2y=5,\]\[3y+2x=10,\]then using matrices and performin gaussian elimination u get\[\begin{pmatrix} 3 & 2 & 5\\ 2 & 3 & 10 \end{pmatrix}\to\begin{pmatrix} 1 & 0 & -1\\ 0 & 1 & 4 \end{pmatrix}\]so \(x=-1\) and \(y=4\)

OpenStudy (anonymous):

hmm what is Gaussian elimination and can i use it during an 8th grade test

OpenStudy (anonymous):

oh u in 8th grade??? nvm this is american college lvl

OpenStudy (anonymous):

oh ha ha

OpenStudy (anonymous):

im 14 they teach us that here in russia

OpenStudy (anonymous):

wow im 14 too

OpenStudy (anonymous):

that is so cool though

OpenStudy (anonymous):

well i have to go but thanks anyway

OpenStudy (anonymous):

yea its a very powerful technique theyre never gonna give you more than 2 equation but u can use it to solve bigger systems like this one\[2x+y-z=8\]\[-3x-y+2z=-11\]\[-2x+y+2z=-3\]\[\begin{pmatrix} 2 & 1 & -1 & 8\\ -3 & 1 & 2 & -11\\ -2 & 1 & 2 & -3 \end{pmatrix}\to\begin{pmatrix} 1 & 0 & 0 & 2\\ 0 & 1 & 0 & 3\\ 0 & 0 & 1 & -1 \end{pmatrix}\]so \(x=2\), \(y=3\) and \(z=-1\)